# Binary population synthesis of the Galactic canonical pulsar population

### v1

@@ 3.2 Results
Figure 4: Top: P-$\dot{P}$ diagram of Fermi detected canonical pulsars (blue) and detected $\gamma$-ray pulsars from the characteristic PN1-B1 model (red); bottom: P-$\dot{P}$ diagram of radio detected canonical pulsars as recorded in the ATNF catalog (blue) and detected radio pulsars from the characteristic PN1-B1 model (red). The two black dashed lines represent the death-lines described in Eqs. 6 &  7. The gray dashed lines labeled $10^{3}$ yr, $10^{6}$ yr, and $10^{9}$ yr represent constant characteristic age each respectively; and those labeled with $10^{31}$ erg s-1 , $10^{33}$ erg s-1 , and $10^{37}$ erg s-1 represent constant $\dot{E}$ each respectively.
In Smith et al. (2019), a discussion of distance biases on the pulsar catalog was presented. Lower $\dot{E}$ pulsars, the majority of which are off the Galactic plane, cannot be detected easily due to their low efficiency, small sample, and possible large distance. However, Smith et al. (2019) suggest that if low $\dot{E}$ pulsars emit in $\gamma$-rays, it would be useful to use the population synthesis approach to help understand if a large population of such pulsars would contribute to a part of the diffuse Galactic $\gamma$-ray emission. This work does not provide a full population modelling of pulsars, especially old pulsars, as only the recent star formation history is considered. Therefore we are unable to determine if the diffuse background can be attributed to old pulsars. This will be discussed in a follow-up study.
Here we show the heuristic flux of detected and stacked pulsars in Fig. 5, as for the detected pulsars, both those from the simulation and the Fermi-LAT catalog occupy similar parts of the parameter space. The heuristic flux is defined as $\sqrt{\dot{E}}/4\pi d^{2}$, and is often used in observation to estimate the potential $\gamma$-ray flux of a pulsar, especially if it is not detectable. In Fig. 5, excluding those with estimated distance 25 kpc (which is the flagged value based on the dispersion measure (Yao et al., 2017)), the stacked pulsars from the simulation, in general, have larger distances compared to those from the catalog. This might also stem from the fact that the distance measurements of the stacked pulsars from the catalog are largely uncertain.
Figure 5: Expected $\gamma$-ray flux of pulsars (the heuristic flux) as a function of the spindown luminosity $\dot{E}$ (similar to Fig. 7 in Smith et al., 2019). We plot the expected $\gamma$-ray flux of pulsars (the heuristic flux) when assuming $L_{\gamma}=\sqrt{10^{33}\dot{E}}$ erg s-1. Blue squares are Fermi-LAT detected pulsars, red triangles are the detected pulsars from the simulation using the PN1-B1 model, green dots are the pulsars from the sample list of Song et al. (2023), and the orange crosses are the pulsars that satisfy the criteria of being stacked from the simulation using the PN1-B1 model. The gray dotted lines labelled as 0.3 kpc, 1 kpc, 5 kpc and 15 kpc represent those of constant distance respectively. Most of the pulsars stacked in Song et al. (2023) with $\dot{E}$ $<10^{33}$ erg s-1 remain undetected in spite of having similar heuristic fluxes to those detected with higher $\dot{E}$. Most of the pulsars being stacked from the simulation have larger distances compared to those from Song et al. (2023).
We show the spatial coordinates of detected pulsars in Figs. 6 &  7. These plots show a similar story as the P-$\dot{P}$ diagram, where the distributions of $\gamma$-ray pulsars across the Galaxy seem to agree better with the catalogs, and the radio pulsars less so. We note the main difference in the spatial distribution of radio pulsars between the catalog and simulation is a large population of radio pulsars around the Galactic Centre is present in the simulation. There is however no such population present in the catalog. There are two possible causes for this issue: 1). The underlying stellar population that produced the pulsars is extremely concentrated around the Galactic Centre, making it the dominant effect. It is reasonable for this to happen, even with the selection effect accounting for high sky temperature around this direction of observation. 2). It is also likely that the distance measurements of the detected radio pulsars are largely uncertain, especially towards the direction of the Galactic Centre, due to high DM/scatter in this direction, causing the catalog pulsars to have inaccurate coordinates.
@@
Figure 6: Galactocentric coordinates of radio pulsars from the catalog (blue) and from PN1-B1 model (red). The yellow star in all panels except for the Y-Z planes indicates the location of the Sun.
@@
Figure 7: Galactocentric coordinates of $\gamma$-ray pulsars from the catalog (blue) and from PN1-B1 model (red). The yellow star in all panels except for the Y-Z planes indicates the location of the Sun.
@@
3.3 Impact of $\gamma$-ray Luminosity Model and Sensitivity
We understand that there are impacts on this work from various model choices implemented in this study. Here we discuss the impact on the simulation results from two model choices: $\gamma$-ray luminosity of pulsars, and the $\gamma$-ray sensitivity limit.
Luminosity determination can impact the simulation results from the uncertainty in the fitting of the fundamental plane itself. As described above in § 2.8, we use the fundamental plane relations with the best-fit parameters without any scatter. Introducing the uncertainties back into the parameters would end up with slightly different fundamental plane relations for each pulsar. We attempt to set the fundamental plane with the uncertainties included in Eq. 17, and re-analyse the PN1-B1 model.
In Fig. 8, we plot the $\gamma$-ray luminosity of the observed pulsars in PN1-B1 model and catalog vs. their $\dot{E}$ for both the original analysis and the $\gamma$-ray luminosity with the scattered fundamental plane. Overall, when taking the fundamental plane without any uncertainties, the Fermi detected pulsars and pulsars from the simulation follow the fundamental plane relationship. The obvious effect of using a scattered fundamental plane is that the luminosity in the model does not follow the apparent fundamental plane relations where $L\sim\dot{E}^{0.39}$. As shown in the right panel of the figure, the correlation is closer to the $L\sim\dot{E}$ line. This is simply due to the fact that the scatter from the power law indices on various quantities have uncertainties that are too big, causing this up scatter. Additionally, some pulsars have luminosity larger than their $\dot{E}$ both in simulation and catalog. When this occurs in the simulation, it is due to the observational effect of the flux correction factor which is caused by the beaming of the pulsar. There are other reasons for this to happen in observed pulsars, most likely due to the large errors in the estimation of pulsar distances and beaming geometry of the pulsars.
Figure 8: Left: $\gamma$-ray luminosity of all detected pulsars vs. $\dot{E}$ , assuming for fixed fundamental plane. Right: $\gamma$-ray luminosity of all detected pulsars vs. their $\dot{E}$ for the scattered fundamental plane. In both panels, the blue data points with error bars are the pulsars detected by Fermi-LAT and the red data points are from the simulation using PN1-B1 model. The magenta dashed line represents the heuristic relation $L_{\gamma}=\sqrt{10^{33}\dot{E}}$ and the red solid line represents the limit of energy conservation as $L_{\gamma}=\dot{E}$.
@@ 3.3 Impact of $\gamma$-ray Luminosity Model and Sensitivity
We also discovered that the choice of the sensitivity limit for the radio-detected pulsars, as described in § 2.8.2, can impact the final results. Specifically, when we change our assumption about the sensitivity floor, the detected population in the model will change too. As stated in § 2.8.2, we deviate from the standard 4FGL sensitivity based on the radio detectability of the pulsar. We define $\alpha$ as the factor to raise the sensitivity floor from the default 4FGL-DR4 sensitivity when the pulsar is not detected in the radio; and $\beta$ as the factor to lower the sensitivity floor when the pulsar is detected in the radio. Here we set a few different options for the combination of $\alpha$ and $\beta$ as described in Table. 6, then re-run the analysis pipelines. We compare the results by checking the number of $\gamma$-ray pulsars produced in each instance, as well as the ratio between radio-quiet and radio-loud $\gamma$-ray pulsars ($q$). This analysis is performed on the PN1-B1 model with 6 different combinations of $\alpha$ and $\beta$. Table. 6 shows the results of the number of $\gamma$-ray pulsars detected in the model (NG) as well as the ratio $q$ between the $\gamma$-ray loud and $\gamma$-ray quiet radio pulsars. When $\alpha$ increases, more $\gamma$-ray pulsars are detected and the ratio between radio-quiet and radio-loud $\gamma$-ray pulsars increases. In this case, most of the increase of NG can be attributed to more radio-quiet pulsars being detected in $\gamma$-rays. When $\beta$ increases, more $\gamma$-ray pulsars are detected and the ratio between radio-quiet and radio-loud $\gamma$-ray pulsars decreases. In this case, most of the increase of NG can be attributed to more radio pulsars being detected in $\gamma$-rays.
Table 6: Results of different setup for the $\gamma$-ray sensitivity. $\alpha$ is the factor to raise the sensitivity floor from the 4FGL value at the location if the pulsar is not detected in radio, and $\beta$ is the factor to lower the sensitivity from the 4FGL value at the location if the pulsar is detected in radio. Model-0 is the same that adopted for the results above. Models 1-6 are additional models. NG is the number of $\gamma$-ray pulsars produced by the model given the sensitivity setup, and the q is the ratio between the radio-quiet and radio-loud $\gamma$-ray pulsars. Cat denotes the observed values for NG and q.
Model $\alpha$ $\beta$ NG q
Cat 160 90:70
0 1.5 5.0 152 87:65
1 1.0 3.0 171 122:49
3 1.0 5.0 171 114:57
3 1.25 4.0 160 99:61
4 2.0 1.0 81 62:19
5 2.0 2.5 98 57:41
6 2.0 4.0 128 68:60
@@
Overall, this further confirms that there are multiple factors that impact the results of this study. Future development in better physical understanding of pulsar emission mechanism and stronger correlation between the radio and $\gamma$-ray detection limit from the current population can further improve these results.
@@ 3.4 Low $\dot{E}$ Pulsars in $\gamma$-rays
Theoretical studies of $\gamma$-ray emission from pulsars, such as the equatorial current sheet model (Pétri, 2012; Kalapotharakos et al., 2018; Hakobyan et al., 2023) or the outer gap model (Cheng et al., 1986a, b; Romani, 1996; Takata et al., 2011), seem to indicate that pulsars with $\dot{E}$ $<10^{33}$ erg/s do not emit in $\gamma$-ray s. Indeed, if we investigate Eq. 17, it shows that at $\dot{E}$ $=10^{33}$ erg/s, the $\gamma$-ray luminosity of a pulsar is equal to $\dot{E}$ , reaching maximum efficiency. With $\dot{E}$ $<10^{33}$, if we still follow the fundamental plane relation, the $\gamma$-ray luminosity of the pulsar becomes larger than $\dot{E}$ , breaking the conservation of energy. It is also worth noting that according to Kalapotharakos et al. (2019, 2022), the fundamental plane relations are rooted in the equatorial current sheets model with theoretical foundations.
However, the most up-to-date 4FGL/3PC catalogs include a handful of $\gamma$-ray pulsars that are detected with $\dot{E}$ $<10^{33}$ erg/s. Given the low number of detections of $\gamma$-ray pulsars with $\dot{E}$ $<10^{33}$ erg/s, there is no clear understanding of their emission mechanism. In fact, according to Song et al. (2023), pulsars might have a universal, weak, and isotropic $\gamma$-ray emission. We try to replicate the stacking analysis performed in Song et al. (2023) to show further evidence of the possible emission mechanism for low $\dot{E}$ pulsars.
@@
We choose pulsars from the PN1-B1 model that are radio loud, $\gamma$-ray quiet that are at least $20^{\circ}$ away from the Galactic plane for this study, consistent with Song et al. (2023). There are 590 such pulsars in the simulation. We then assign each pulsar with two different $\gamma$-ray luminosities, one that follows the fundamental plane if $\dot{E}>10^{33}$ erg s-1and $L=0.8\dot{E}$ when $\dot{E}<10^{33}$ erg s-1; and the other one is assigned a luminosity drawn from a log normal distribution $\log_{10}\sim N(32,0.5)$ (ergs/s). The former luminosity assignment corresponds to beamed $\gamma$-ray emission correlated to the $\dot{E}$ of each pulsar, and the latter corresponds to the universal, weak and isotropic $\gamma$-ray emission. For the first emission mechanism, we still account for the $\gamma$-ray beaming effect discussed in § 2.8.3 when being stacked, while the isotropic emission does not need to be beamed.
@@
We then calculate the average luminosity of the pulsars in five different $\dot{E}$ bins, with each bin having the same number of pulsars. We check these results in a similar way to Fig. 6 in Song et al. (2023). Here in this work, we choose the canonical pulsars from Song et al. (2023) and plot them in Fig. 9, and use this as the observational results to be compared with the model. We then plot in Fig. 9 the average luminosity vs. $\dot{E}$ from the model with three different scenarios: only the fundamental plane is considered, only isotropic emission is considered, and combining both. Unfortunately, because the observational results are largely uncertain at this point, we are unable to determine which would be a better fit to describe the emission mechanism of lower $\dot{E}$ pulsars. We either need a better understanding of pulsar emission in this regime, or more data to improve the observational results.
@@
Figure 9: Left: $\gamma$-ray luminosity vs. $\dot{E}$ for stacked pulsars. The green data points with error bars in both directions are from Song et al. (2023) and the orange data points with only horizontal error bars are from the simulation using PN1-B1 model. The magenta dashed line represents the heuristic relation $L_{\gamma}=\sqrt{10^{33}\dot{E}}$ and the red solid line represents the limit of energy conservation as $L_{\gamma}=\dot{E}$. The left panel calculates the stacked luminosity from the simulation by assuming the stacked pulsars have a luminosity that follows the fundamental plane while $\dot{E}>10^{33}$ erg s-1 and $L_{\gamma}=0.8\dot{E}$ when $\dot{E}<10^{33}$ erg s-1Ṫhe middle panel is calculating the luminosity assuming all these pulsars have a weak, isotropic $\gamma$-ray emission follows a log-normal distribution where $\log{L_{\gamma}}~{}N(32,0.5)$. The right panel combines both.
@@
Another way to examine the results and potential emission mechanism is to compare the spectral energy distributions (SEDs) for various scenarios. We calculate the stacked SED of the 590 pulsars for the fundamental plane emission and isotropic emission as described above. The $\gamma$-ray spectral parameters of the pulsars, $E_{\textrm{cut}}$ and PL index before cutoff, are drawn from the described text above. When considering the fundamental plane emission, a pulsar needs to be beamed towards the observer. When considering the isotropic emission, it does not need to be beamed. Averaging the SEDs, we can produce the stacked SEDs for both scenarios as shown in Fig. 10. We compare this result to Song et al. (2023) as well as the averaged pulsar SED from McCann (2015). As expected, if considering the FP emission, this stacked pulsar population is relatively older compared to the young, energetic pulsars, hence it is somewhere in between the averaged young pulsar and MSP SED. That does mean that this is much more luminous than the stacked pulsars from Song et al. (2023). When considering the isotropic emission, however, the SED is much less luminous compared to the stacked pulsars from the data. Neither of our models here can completely explain the observation. It might be a mixture of both scenarios that can account for the observation, averaging out to the observed data points.
@@
Figure 10: SED of stacked pulsars from the simulation using the PN1-B1 model. The orange solid curve is the averaged SED from the stacked pulsars in the simulation assuming they all follow the empirical fundamental plane (FP) emission mechanism in $\gamma$-rays, and the red dashed curve is for the same population of pulsars assuming they all emit weakly isotropic $\gamma$-ray emission (ISO). The blue data points with error bars are the averaged SED of the sampled canonical pulsars from Song et al. (2023), The green dotted-dashed curve represents the averaged SED of young pulsars, and the green dotted curve SED of MSPs (McCann, 2015).
@@
Many low $\dot{E}$ pulsars are located outside of the Galactic plane (Smith et al., 2023). It is commonly believed that the high latitude pulsars are a result of pulsars receiving larger kicks when they went through supernova explosions. The results from this work can be used to confirm if that is indeed the case, as the kick velocities are one of the modelled components in our population synthesis models. In Fig. 11, we show the distributions of kick velocities received by pulsars from the model that are either detected or stacked, as selected above. Consistent with our hypothesis, we find that the high latitude, low $\dot{E}$ pulsars in the stacked sample received larger natal kicks than the typical detected population.

### v2

@@ 3.5 Best Fit Model
Figure 4: Top: P-$\dot{P}$ diagram of Fermi detected canonical pulsars (blue) and detected $\gamma$-ray pulsars from the characteristic PN1-B1-D4 model (red); bottom: P-$\dot{P}$ diagram of radio detected canonical pulsars as recorded in the ATNF catalogue (blue) and detected radio pulsars from the characteristic PN1-B1-D4 model (red). The two black dashed lines represent the death-lines described in Eqs. 7 &  8. The gray dashed lines labelled $10^{3}$ yr, $10^{6}$ yr, and $10^{9}$ yr represent constant characteristic age each respectively; and those labelled with $10^{31}$ erg s-1 , $10^{33}$ erg s-1 , and $10^{37}$ erg s-1 represent constant $\dot{E}$ each respectively.
We show the spatial coordinates of detected pulsars in Figures 5 &  6. In both Figures 5 &  6, the spatial distributions of pulsars in the X-Y, X-Z and Y-X planes have good overlap between the observation and the simulation, with a slight over density around the Galactic Centre region. This over density is the most obvious in the R-Z plot at the bottom right panel in both Figures 5 &  6. We attribute this to the initial spatial distributions of massive stars, but overall the Galactic model used in this work is sufficiently good.
@@
Figure 5: Galactocentric coordinates of radio pulsars from the catalogue (blue) and from PN1-B1-D4 model (red). The yellow star in all panels except for the Y-Z planes indicates the location of the Sun.
@@
Figure 6: Galactocentric coordinates of $\gamma$-ray pulsars from the catalogue (blue) and from PN1-B1-D4 model (red). The yellow star in all panels except for the Y-Z planes indicates the location of the Sun.
3.6 Magnetic Decay Timescale
Magnetic decay timescale ($\tau_{d}$) is a key parameter for pulsar evolution models. As mentioned in § 2.4, there is no consensus on the value of $\tau_{d}$. Some studies, mentioned in § 2.4, suggest short $\tau_{d}$, indicating rapid magnetic field decay; while others prefer long $\tau_{d}$, indicating a slow-changing or even constant magnetic field throughout the life time of a NS. In this section, we use population synthesis methods to examine both perspectives. As shown in Table 1, four sets of models are examined: INIT, P2, PN1-B1 and PN5-B5, each with specific birth magnetic field and spin period distributions and $\tau_{d}$ ranging from 1 Myr to 10,000 Myr.
For radio pulsars, shorter $\tau_{d}$ values (1–10 Myr) generally produce better matches to observations across most model sets, while longer $\tau_{d}$ values (100–10,000 Myr) tend to yield worse fits. The caveat of having shorter $\tau_{d}$ for radio pulsars is that the models tend to generate more low field pulsars, resulting with a population with shorter spin periods. For $\gamma$-ray pulsars, moderate to long $\tau_{d}$ values (e.g., ¿ 100 Myr) better match the observed population, as slow-decaying or effectively non-decaying magnetic fields are more consistent with the high $\dot{E}$ s and low characteristic ages of these pulsars. Notably, model set PN1-B1 stands out for its strong performance in describing radio pulsar populations, with $\tau_{d}=1$ Myr emerging as the best fit.
Overall, the results favour shorter $\tau_{d}$ values (1–10 Myr) for describing the population of neutron stars, particularly radio pulsars, as these models consistently align with observed distributions of pulsar properties, as well as the numbers and ratios of different pulsar types. However, $\gamma$-ray pulsars require longer $\tau_{d}$ values for optimal fits.
One caveat is that this result is based on the assumption that the magnetic field always decays exponentially with time, as described in § 2.3. The exponentially decaying field in our model suggests that different decay timescales are needed to explain older radio and younger gamma-ray pulsars. However, accounting for the Hall effect at earlier times, as mentioned in § 2.3, would bridge this issue. This would require modelling the magnetic field decay with different prescriptions (Colpi:2000ApJL; DallOsso:2012MNRAS, e.g.,) and it is planned for future work (see § 4).
3.7 Impact of $\gamma$-ray Beaming and Sensitivity
@@
Two model choices can impact the simulation results: $\gamma$-ray beaming and the $\gamma$-ray sensitivity limit.
Our initial suspicion for the underproduction of $\gamma$-ray pulsars was that it was due to the beaming factor (as described in § 2.8.3). Since $\gamma$-ray beaming is applied as a probabilistic rejection scheme, variations in the beaming will impact the numbers of $\gamma$-ray pulsars. We made modifications to the beaming prescription outlined as follows: Beaming 1 option is the one described in § 2.8.3. Beaming 2 assumes the beaming fraction remains high to a floor of fg = 0.7 for all pulsars with $\dot{E}$ ¡ $10^{35}$ erg s-1. For pulsars with $10^{35}$ erg s-1 ¡ $\dot{E}$ ¡ $10^{36}$ erg s-1, fg = 0.75. For pulsars with $10^{36}$ erg s-1 ¡ $\dot{E}$ ¡ $10^{37}$ erg s-1, fg = 0.8. Beaming 3 decreases the beaming fraction of pulsars by steps in the same way as Beaming 1 until a floor at fg = 0.4 at $\dot{E}$ = 1035 erg s-1. These beaming models are simply ad hoc generalisation made to compare different prescriptions. All three beaming options are plotted in Figure 7.
Figure 7: Three models for the beaming of $\gamma$-ray pulsars as a function of $\dot{E}$ as described in Section 3.7.
@@
The sensitivity limit for radio-detected $\gamma$-ray pulsars, as described in § 2.8.2, can also affect the final results. Specifically, when we change our assumption about the sensitivity limit, the detected population in the model will also change. As stated in § 2.8.2, we deviate from the standard 4FGL sensitivity based on the radio detectability of the $\gamma$-ray pulsar. Here we explore the effects of varying $1.0\leq\alpha\leq 4.0$ and $1.0\leq\beta\leq 4.0$, then re-run the analysis pipelines with all three beaming options on the PN1-B1-D4 model. We compare the resulting number of $\gamma$-ray pulsars produced in each instance (NG) and the ratio of $\gamma$-ray pulsars and radio-loud $\gamma$-ray pulsars ($q$) to the catalogue values in Figure 8.
Figure 8: Results showing NG (yellow-brown contours) and q (blue-green contours) with different combination of $\alpha$ and $\beta$, under Beaming 1 (left), Beaming 2 (middle) and Beaming 3 (right) options, respectively. The catalogue values of NG = 41 and q = 2.24 are labelled as dashed lines in the contours. The fuchsia cross in each panel represents the default choices of $\alpha=3.0$ and $\beta=1.5$ used in the analysis as stated in § 2.8.2. The red star in each panel denotes the intersection where the simulation results match both NG and q.
The results of this analysis are plotted in Figure 8 with contours of NG and q for different combinations of $\alpha$ and $\beta$ for all three beaming options. Assuming the catalogue contains the same number of radio pulsars as predicted by the PN1-B1-D4 model (509; see Table 5), the number of catalogue $\gamma$-ray pulsars is scaled down proportionally from 85 to 41, reflecting the same reduction ratio as applied to the radio pulsars (from 1060 to 509), to ensure a fair comparison. NG = 41, and the catalogue value of q = 2.24 are both plotted with dashed lines in the contours. Firstly, the impact of beaming is examined. Beaming 2 increases number of $\gamma$-ray pulsars for all combinations of $\alpha$ and $\beta$, whilst showing a slight decrease in q in most cases. Beaming 3 shows marginal difference from Beaming 1.
For a given beaming option, when $\alpha$ increases, fewer $\gamma$-ray pulsars are detected, and q decreases. In this case, the decrease of NG can be attributed to radio-quiet pulsars being harder to detect in $\gamma$-rays without existing timing solutions. When $\beta$ increases, more $\gamma$-ray pulsars are detected, but q decreases. This can be attributed to more radio pulsars being detected in $\gamma$-rays. In all three different beaming options, high $\alpha\gtrsim 2.8$ is required to achieve the catalogue values of q and NG. This agrees with the initial assumption that according to the 2PC/3PC catalogues Fermi-LAT:2013svs; Fermi-LAT:2023zzt, Fermi-LAT is less sensitive to pulsars discovered without known radio counterparts (e.g. through blind searches). In beaming options 1 and 3, high $\beta$ is required to achieve the catalogue values, which suggests that pulsars with prior radio detection is much easier to be detected in Fermi-LAT. In beaming option 2, while only a low $\beta$ is required, the beaming fractions for pulsars with different $\dot{E}$ are all raised significantly. This is another indication that radio pulsars should be easier to detect in $\gamma$-ray to agree with the catalogues. In comparison, if choosing the unaltered Fermi-LAT sensitivity with $\alpha=\beta=1$, the models overproduce $\gamma$-ray pulsars, and also produce more radio quiet pulsars indicated by higher q.
@@
Future development in better physical understanding of $\gamma$-ray beaming is required. It is also important to establish a correlation between the radio and $\gamma$-ray detection limit from the current population, which can further improve results for future population study.
@@ 3.8 Low $\dot{E}$ Pulsars in $\gamma$-rays
Theoretical studies of $\gamma$-ray emission from pulsars, such as the equatorial current sheet model current_sheet0; current_sheet1; current_sheet2 or the outer gap model outer_cheng1986a; outer_cheng1986b; outer_romani1996; Takata2011, seem to indicate that pulsars with $\dot{E}$ $10^{33}$ erg/s do not emit in $\gamma$-ray s. Indeed, if we investigate Eq. 17, it shows that at $\dot{E}$ $10^{33}$ erg/s, the $\gamma$-ray luminosity of a pulsar is equal to $\dot{E}$ , reaching maximum efficiency. With $\dot{E}$ $10^{33}$ erg s-1 , if we still follow the fundamental plane relation, the $\gamma$-ray luminosity of the pulsar becomes larger than $\dot{E}$ , breaking conservation of energy. It is also worth noting that according to Pulsar_FP2019; Pulsar_FP2022, the fundamental plane relations are rooted in the equatorial current sheets model with theoretical foundations.
[…]
@@
We choose pulsars from the PN1-B1-D4 model that are radio-loud, $\gamma$-ray quiet, that are at least $20^{\circ}$ away from the Galactic plane for this study, consistent with Song2023. There are 46 such pulsars in the randomly chosen realisation of the simulation. We then assign each pulsar with two different $\gamma$-ray luminosities, one that follows the fundamental plane if $\dot{E}>10^{33}$ erg s-1and $L=0.8\dot{E}$ when $\dot{E}<10^{33}$ erg s-1; and the other one is assigned a luminosity drawn from a log normal distribution $\log_{10}\sim N(32,0.5)$ (ergs/s). The former luminosity assignment corresponds to beamed $\gamma$-ray emission correlated to the $\dot{E}$ of each pulsar, and the latter corresponds to the universal, weak and isotropic $\gamma$-ray emission. For the first emission mechanism, we still account for the $\gamma$-ray beaming effect discussed in § 2.8.3 when being stacked, while the isotropic emission does not need to be beamed.
@@
We then calculate the average luminosity of the pulsars in four different $\dot{E}$ bins, with the lowest two $\dot{E}$ bins have the same number of pulsars and highest two $\dot{E}$ bins having the same number of pulsars. We check these results in a similar way to Figure 6 in Song2023. Here in this work, we choose the canonical pulsars from Song2023 and plot them in Figure 9, and use this as the observational results to be compared with the model. We then plot in Figure 9 the average luminosity vs. $\dot{E}$ from the model with three different scenarios: only the fundamental plane is considered, only isotropic emission is considered, and combining both. While the observational results are largely uncertain, they are consistent with all three scenarios proposed in this work.
@@
Figure 9: $\gamma$-ray luminosity vs. $\dot{E}$ for stacked pulsars. The green data points with error bars in both directions are from Song2023 and the orange data points with only horizontal error bars are from the simulation using PN1-B1 model. The magenta dashed line represents the heuristic relation $L_{\gamma}=\sqrt{10^{33}\dot{E}}$ and the red solid line represents the limit of energy conservation as $L_{\gamma}=\dot{E}$. The left panel calculates the stacked luminosity from the simulation by assuming the stacked pulsars have a luminosity that follows the fundamental plane while $\dot{E}>10^{33}$ erg s-1 and $L_{\gamma}=0.8\dot{E}$ when $\dot{E}<10^{33}$ erg s-1Ṫhe middle panel is calculating the luminosity assuming all these pulsars have a weak, isotropic $\gamma$-ray emission follows a log-normal distribution where $\log{L_{\gamma}}~N(32,0.5)$. The right panel combines both.
@@
We further examine the results and potential emission mechanism is to compare the spectral energy distributions (SEDs) for various scenarios. We calculate the stacked SED of the 46 pulsars for the fundamental plane emission and isotropic emission as described above. The $\gamma$-ray spectral parameters of the pulsars, $E_{\textrm{cut}}$ and PL index before cutoff, are drawn from the described text above. When considering the fundamental plane emission, a pulsar needs to be beamed towards the observer. When considering the isotropic emission, it does not need to be beamed. Averaging the SEDs, we can produce the stacked SEDs for both scenarios as shown in Figure 10. We compare this result to Song2023 as well as the averaged pulsar SED from 2015ApJ...804...86M. If considering the FP emission, this stacked pulsar population is relatively older compared to the young, energetic pulsars, hence the SED is much lower. Interestingly, when considering the isotropic emission, the SED is more luminous compared to the FP emission, and roughly agrees with the stacked pulsars from the data. By examining the stacked SEDs of modelled pulsars, it seems to indicate that the isotropic emission is favoured over the fundamental plane emission. However the caveat here is that the model produces a lot fewer pulsars that satisfy the stacking criteria.
@@
Figure 10: SED of stacked pulsars from the simulation using the PN1-B1-D4 model. The orange solid curve is the averaged SED from the stacked pulsars in the simulation assuming they all follow the empirical fundamental plane emission description in $\gamma$-rays, and the red dashed curve is for the same population of pulsars assuming they all emit weakly isotropic $\gamma$-ray emission (ISO). The blue data points with error bars are the averaged SED of the sampled canonical pulsars from Song2023, The green dotted-dashed curve represents the averaged SED of young pulsars, and the green dotted curve SED of MSPs 2015ApJ...804...86M.
@@
Many low $\dot{E}$ pulsars are located outside of the Galactic plane Fermi-LAT:2023zzt. It is commonly believed that the high latitude pulsars are a result of pulsars receiving larger kicks when they went through supernova explosions. The results from this work can be used to confirm if that is indeed the case, as the kick velocities are one of the modelled components in our population synthesis models. In Figure 11, we show the distributions of kick velocities received by pulsars from the model that are either detected or stacked, as selected above. Consistent with our hypothesis, we find that the high latitude, low $\dot{E}$ pulsars in the stacked sample received larger natal kicks than the typical detected population.
@@
In Smith:2019ApJ, a discussion of distance biases on the pulsar catalogue was presented. Lower $\dot{E}$ pulsars, the majority of which are off the Galactic plane, cannot be detected easily due to their low efficiency, small sample, and possible large distance. However, Smith:2019ApJ suggest that if low $\dot{E}$ pulsars emit in $\gamma$-rays, it would be useful to use the population synthesis approach to help understand if a large population of such pulsars would contribute to a part of the diffuse Galactic $\gamma$-ray emission. Diffuse Galactic emission has been observed in GeV energies by Fermi-LAT Abdo:2009ApJ and TeV energies (Abramowski:2014PhRvD, e.g. by HESS;), with its origin still under research diMauro:arXiv16 and unresolved pulsars being potential contributors Smith:2019ApJ. This work does not provide a full population modelling of pulsars, especially old pulsars, as only the recent star formation history is considered. Therefore we are unable to determine if the diffuse background can be attributed to old pulsars. This will be discussed in a follow-up study.
Here we show the heuristic $\gamma$-ray flux of detected and stacked pulsars in Figure 12, as for the detected pulsars, both those from the simulation and the Fermi-LAT catalogue occupy similar parts of the parameter space. The heuristic $\gamma$-ray flux is defined as $\sqrt{\dot{E}}/4\pi d^{2}$, and is often used in observation to estimate the potential $\gamma$-ray flux of a pulsar, especially if it is not detectable. We exclude pulsars that are estimated to have distances equal to 25 kpc, where the distances are estimated based on the dispersion measure of radio observations, and rely on an underlying electron density model YMW17. In Figure 12, the stacked pulsars from the simulation, in general, occupy similar ranges of distances compared to those from the catalogue. However the distribution of the distances from the stacked pulsars in the simulation peaks at larger distance compared to those from the catalogue.
Figure 12: Expected $\gamma$-ray flux of pulsars (the heuristic flux) as a function of the spindown luminosity $\dot{E}$ (similar to Figure 7 in Smith:2019ApJ). We plot the expected $\gamma$-ray flux of pulsars (the heuristic flux) when assuming $L_{\gamma}=\sqrt{10^{33}\dot{E}}$ erg s-1. Blue squares are Fermi-LAT detected pulsars, red triangles are the detected pulsars from the simulation using the PN1-B1 model, green dots are the pulsars from the sample list of Song2023, and the orange crosses are the pulsars that satisfy the criteria of being stacked from the simulation using the PN1-B1 model. The gray dotted lines labelled as 0.3 kpc, 1 kpc, 5 kpc and 15 kpc represent those of constant distance respectively. Most of the pulsars stacked in Song2023 with $\dot{E}$ $10^{33}$ erg s-1 remain undetected in spite of having similar heuristic fluxes to those detected with higher $\dot{E}$. Most of the pulsars being stacked from the simulation have larger distances compared to those from Song2023.